Notation for the Probability of the given event:

1) What percentage satisfied exactly one of the three criteria?
2) Of those who satisfied at least two of the three criteria, what percentage satisfied exactly two criteria?
Answer #1

1)

What percentage satisfied exactly one of the three criteria?

Total = 1 + 4 + 5 + 7 + 2 + 6 + 3 = 28

Favorable = exactly 1 = 1 + 2 + 3 = 6

Required percentage = 6/28 = 21.4285714285%

2)

Of those who satisfied at least two of the three criteria, what percentage satisfied exactly two criteria?

Total = at least two = 4 + 5 + 7 + 6 = 22

Favorable = exactly 2 = 4 + 5 + 6 = 15

Required percentage = 15/22 = 68.1818181818%